# De-excitations of highly excited 11B∗ and 15N∗ based on the GEMINI++ code

### v1

@@ I Introduction
Excited nuclei can be created by neutrino-nucleon interactions and nucleon decays in neutrino experiments. Subsequent de-excitation processes will emit gamma rays and other massive particles, such as neutrons, protons, deuterons, tritons, and particles. Detecting emitted particles is very helpful for neutrino experiments to analyze some physical topics. Water Cherenkov (WC) detectors had utilized de-excitation gamma rays to measure the nucleon decay modes of [1, 2] and neutron invisible decays [2, 3, 4]. Neutrons can be easily identified in liquid scintillator (LS) experiments based on the delayed signal from neutron capture. LS detectors had applied neutron multiplicity to search for [5, 6], the diffuse supernova neutrino background [7], the strange axial coupling constant [8, 9] and neutron invisible decays [10, 11, 12]. Liquid Argon detectors have the capability to detect all particles from nuclear de-excitations, which can be used to improve the energy resolution of accelerator neutrinos [13] and supernova neutrinos [14]. As new-generation neutrino experiments such as JUNO[15], Hyper-Kamiokande[16], and DUNE[17] will begin collecting data in the near future, nuclear de-excitations will play a more and more significant role.
Accurate de-excitation branching ratios are crucial for neutrino experiments to use the information of de-excitation products. However, there are no universally adopted and quantitatively accurate models to describe the full de-excitation chains [18]. Some theoretical studies focused on the emission associated with one [2] and two [4] nucleon decays in 16O. Experimental measurements of nuclear de-excitations can provide useful and indicative results [19, 20, 22, 21]. However, it is very challenging to measure the complete branching ratios. Consequently, some statistical model codes are applied to calculate de-excitation processes, including TALYS [23] used in Refs.[5, 6, 7, 8, 9, 14], ABLA [24] in Ref.[13, 25], SMOKER [26] in Refs.[12] and CASCADE [27] in Refs.[19, 20]. Note that none of these codes can adequately account for the measured de-excitation results of light nuclei and [19, 20, 6, 9, 25]. In addition, the output of widely used TALYS is not in an event-by-event format, making it difficult to be integrated into neutrino generators, such as GENIE [28]. Therefore, current neutrino experiments urgently require a de-excitation code that can well describe experimental data and provide de-excitation results on an event-by-event basis.
@@ II De-excitations of highly excited 11B∗ and 15N∗
Yosoi et al.[19, 20] measured charged particle emissions from the -hole states of in coincidence with the quasifree reaction using a 392 MeV proton beam, where the charged particles include protons, deuterons, tritons and particles. The detection thresholds are 3.1, 4.0, 4.6 and 4.5 MeV for and emissions, respectively. Note that events have been included in the portion. Due to the detector limitation, the residual daughter nucleus after a full de-excitation cascade can not be detected. In fact, the excitation energy (res) of residual daughter nuclei can be induced from the excitation energy of and the kinetic energy of a detected charged particle as shown in Fig. 2 of Ref. [19]. Then the authors define the ‘two-body decay’ and ‘three-body decay’ in terms of (res). The measured branching ratios of ‘two-body decay’ and ‘three-body decay’ for 16-35 MeV have been plotted in the left panel of Fig. 1. The hatched areas indicate the ‘two-body decay’ regions, and the blank parts mainly come from contributions of three-body and sequential de-excitation processes. It is clear that the and decays contribute the most to the charged particle emissions from the -hole states of excited . Most of , and particles belong to the ‘two-body decay’.
Yosoi et al.[20] also measured charged particle emissions from de-excitations of produced by the quasifree reaction using the same proton beam and detector. The right panel of Fig. 1 shows the measured branching ratios of charged particles and neutrons from de-excitations for 20-40 MeV. The fraction of and decays is smaller than that of and decays, which is completely different from the case. It is worthwhile to stress that this experiment measured separately the neutron emission of highly excited in the second beam time [20]. To detect neutrons, this measurement used a small scattering chamber and a neutron multi-detector array system. The lower limit of neutron detection energy is 3.2 MeV [32]. As shown in the right panel of Fig. 1, the measured branching ratio of neutrons is about 50%, which is larger than the contribution of all charged particles. For the neutron emission, the contributions of both the ‘two-body decay’ and ‘three-body decay’ are roughly equal.
@@ III GEMINI++ predictions
where and are the spin and orbital angular momenta of the evaporated particle , respectively. Evaporation modes include the emissions of , , , , , , and so on. is the transmission coefficient or barrier penetration factor, and depends on the kinetic energy of the evaporated particle. in Eq. (1) represents , where is the separation energy and is the rotational energy of the ground-state configuration [36]. and are the spin-independent level densities of the parent and daughter nuclei, and are often approximated by the Fermi-gas form [31]:
@@ IV.1 Modifying GEMINI++
As shown in Eq. (2), the effect of the back-shifted term is equivalent to reducing the excitation energy of the compound nucleus. It can significantly suppress the emissions of massive particles compared with the emission. Therefore, massive particles will necessitate a higher excitation energy to initiate their evaporation process. To determine this critical excitation energy for 11B∗ de-excitations in the GEMINI++ code, we simulate 5000 events for every interval of 0.1 MeV in the range of . The critical energies for 5 two-body decay modes have been listed in Table. 1. The predicted for decay modes of and exhibit obvious deviations from their separation energies. There should not be such large differences though additional energy is needed to overcome the nuclear effects. This implies that the back-shifted term is not used properly for the case of light nuclei. Compared with the fitting results in Ref. [39], used in GEMINI++ indicates a significant overestimation due to the large pairing correction calculation. Unfortunately, the available is absent for 11B∗ and 15N∗. Therefore, we modify GEMINI++ through removing the back-shifted term in Eq. (2). In this case, the critical energy is closest to the separation energy. As a result, removing can enhance the and branching ratios, and correspondingly reduce the and emissions. For 11B∗, this modification will bring the predictions closer to the experimental observations. However, the predicted branching ratios of , and from 15N∗ de-excitations will deviate from their measured values.
@@
The daughter nuclei of de-excitations are usually left in discrete levels when they have a relatively low excitation energy. However, this situation does not appear in GEMINI++ since it only considers continuum levels. Here we calculate the influence of discrete levels, which can be implemented through modifying the GEMINI++ outputs on an event-by-event basis. Due to the unclear boundary between discrete and continuum states, we only consider the discrete levels with the excitation energy less than about 6 MeV, and require that the decay information of every added discrete state is known clearly in the NNDC database [40]. Modifications of the GEMINI++ output events can be categorized into three types of situations:
If the excited energy of parent nucleus in the last binary decay is higher than the highest discrete level considered obviously, this GEMINI++ event will be kept;
@@
Based on the above strategy, we modify the GEMINI++ outputs and find that including discrete levels can change the kinetic energy distributions of emitted particles. These variations can affect the predicted branching ratios of both ‘two-body decay‘ and ‘three-body decay’ in Fig. 1. Note that the setting of discrete levels does not change the predicted relative branching ratios in Fig. 2.
@@ IV.2 GEMINI++4 results
Here we discuss the GEMINI++4 performance compared with the experimental data in Refs. [19, 20, 21]. Both and scenarios have removed the back-shifted term and added the discrete levels. For the setting, the predicted branching ratios of 11B∗ de-excitations are in a good agreement with the measurements as shown in the left panel of Fig. 1. This is the first time that a statistical model code can simultaneously explain all four emissions of charged particles for both two-body and three-body decays of 11B∗, although there are still some minor differences. However, GEMINI++4 with can not account for the 15N∗ de-excitations well. It is clear that obvious differences in the right panel of Fig. 1 can be found for the , and decays. For the relative branching ratios in Fig. 2, GEMINI++4 with gives the better predictions than those from TALYS [6], NucDeEx [9] and ABLA [25].
@@ V Summary
In summary, we have investigated the de-excitations of highly excited 11B∗ and 15N∗ by use of the GEMINI++ code. Due to three serious issues of the critical excitation energy, discrete levels and suppression factors, GEMINI++ can not account for the experimental data well. After addressing the three issues, we develop a code of GEMINI++4 for neutrino experiments to handle de-excitaions of light nuclei based on the GEMINI++ code. For the settings of suppression factors, and have been assumed in this work. The predicted branching ratios of 11B∗ de-excitations are consistent with the measurements in the case. However, this setting can not explain the 15N∗ data. GEMINI++4 with can give the better agreement with the 11B∗ data, and partially account for the measured branching rations of 15N∗ de-excitations. Therefore, the setting is recommended for the relevant applications. This is the first time that a statistical model code can basically reproduce both 11B∗ and 15N∗ data. Moreover, the comparisons for each energy bin show that the agreement between the GEMINI++4 predictions and the experimental data is not coincidental.
@@ Acknowledgements.
- [38] A. J. Koning, S. Hilaire and S. Goriely, Nucl. Phys. A 810 (2008), 13-76 doi:10.1016/j.nuclphysa.2008.06.005

### v2

@@ I Introduction
Excited nuclei can be created by neutrino-nucleon interactions and nucleon decays in neutrino experiments. Subsequent de-excitation processes will emit gamma rays and other massive particles, such as neutrons, protons, deuterons, tritons, and particles. Detecting emitted particles is very helpful for neutrino experiments to analyze some physical topics. Water Cherenkov (WC) detectors had utilized de-excitation gamma rays to measure the nucleon decay modes of [1, 2] and neutron invisible decays [2, 3, 4]. Neutrons can be easily identified in liquid scintillator (LS) experiments based on the delayed signal from neutron capture. LS detectors had applied neutron multiplicity to search for [5, 6], the diffuse supernova neutrino background [7], the strange axial coupling constant [8, 9] and neutron invisible decays [10, 11, 12]. Liquid Argon detectors have the capability to detect all particles from nuclear de-excitations [13][14], which can potentially be used to improve the energy resolution of accelerator neutrinos [15], supernova neutrinos [16] and other possible physics applications [17]. As new-generation neutrino experiments such as JUNO[18], Hyper-Kamiokande[19], and DUNE[20] will begin collecting data in the near future, nuclear de-excitations will play a more and more significant role.
Accurate de-excitation branching ratios are crucial for neutrino experiments to use the information of de-excitation products. However, there are no universally adopted and quantitatively accurate models to describe the full de-excitation chains [21]. Some theoretical studies focused on the emission associated with one [2] and two [4] nucleon decays in 16O. Experimental measurements of nuclear de-excitations can provide useful and indicative results [22, 23, 25, 24]. However, it is very challenging to measure the complete branching ratios. Consequently, some statistical model codes are applied to calculate de-excitation processes, including TALYS [26] used in Refs.[5, 6, 7, 8, 9, 16], ABLA [27] in Ref.[15, 28], SMOKER [29] in Refs.[12] and CASCADE [30] in Refs.[22, 23]. Note that none of these codes can adequately account for the measured de-excitation results of light nuclei and [22, 23, 6, 9, 28]. In addition, the output of widely used TALYS is not in an event-by-event format, making it difficult to be integrated into neutrino generators, such as GENIE [31] and NuWro [32]. Therefore, current neutrino experiments urgently require a de-excitation code that can well describe experimental data and provide de-excitation results on an event-by-event basis.
@@ II De-excitations of highly excited 11B∗ and 15N∗
Yosoi et al.[22, 23] measured charged particle emissions from the -hole states of in coincidence with the quasifree reaction using a 392 MeV proton beam, where the charged particles include protons, deuterons, tritons and particles. The detection thresholds are 3.1, 4.0, 4.6 and 4.5 MeV for and emissions, respectively. Note that events have been included in the portion. Due to the detector limitation, the residual daughter nucleus after a full de-excitation cascade can not be detected. In fact, the excitation energy (res) of residual daughter nuclei can be induced from the excitation energy of and the kinetic energy of a detected charged particle as shown in Fig. 2 of Ref. [22]. Then the authors define the ‘two-body decay’ and ‘three-body decay’ in terms of (res). The ‘two-body decay’ region for each particle is defined as (res) Max(5 MeV, (3-body)), where (3-body) denotes the threshold energy of massive particle decay in the residual nucleus. The ‘three-body decay’ mainly comes from contributions of the realistic three-body and sequential binary de-excitation processes. Note that ‘two-body decay’ and ‘three-body decay’ are enclosed in single quotes to match the measurements in Refs. [22, 23]. The measured branching ratios of ‘two-body decay’ and ‘three-body decay’ for 16-35 MeV have been plotted in the left panel of Fig. 1. The hatched/colored and blank areas indicate the ‘two-body decay’ and ‘three-body decay’ contributions, respectively. It is clear that the and decays contribute the most to the charged particle emissions from the -hole states of excited . Most of , and particles belong to the ‘two-body decay’.
Yosoi et al.[23] also measured charged particle emissions from de-excitations of produced by the quasifree reaction using the same proton beam and detector. The right panel of Fig. 1 shows the measured branching ratios of charged particles and neutrons from de-excitations for 20-40 MeV. The fraction of and decays is smaller than that of and decays, which is completely different from the case. It is worthwhile to stress that this experiment measured separately the neutron emission of highly excited in the second beam time [23]. To detect neutrons, this measurement used a small scattering chamber and a neutron multi-detector array system. The lower limit of neutron detection energy is 3.2 MeV [36]. Definitions of ‘two-body decay’ and ‘three-body decay’ for are similar with but the maximum value is chosen from 8 MeV and the corresponding (3-body). As shown in the right panel of Fig. 1, the measured branching ratio of neutrons is about 50%, which is larger than the contribution of all charged particles. For the neutron emission, the contributions of both the ‘two-body decay’ and ‘three-body decay’ are roughly equal.
@@ III GEMINI++ predictions
where and are the spin and orbital angular momenta of the evaporated particle , respectively. Evaporation modes include the emissions of , , , , , , and so on. is the transmission coefficient or barrier penetration factor, and depends on the kinetic energy of the evaporated particle. To ensure full absorption, the incoming-wave boundary-condition model (IWBC) [40] is employed to calculate with the help of global optical-model potentials [34]. in Eq. (1) represents , where is the separation energy and is the rotational energy of the ground-state configuration [41]. and are the spin-independent level densities of the parent and daughter nuclei, and are often approximated by the Fermi-gas form [35]:
@@ IV.1 Modifying GEMINI++
As shown in Eq. (2), the effect of the back-shifted term is equivalent to changing the excitation energy of the compound nucleus. It can significantly change the emissions of massive particles compared with the emission. Therefore, massive particles will necessitate a higher excitation energy to initiate their evaporation process. To determine this critical excitation energy for 11B∗ de-excitations in the GEMINI++ code, we simulate 5000 events for every interval of 0.1 MeV in the range of . The critical energies for 6 two-body decay modes have been listed in Table 1. The predicted for decay modes of and exhibit obvious deviations from their separation energies. There should not be such large differences though additional energy is needed to overcome the nuclear effects. This implies that the back-shifted term is not used properly for the case of light nuclei. Compared with the fitting results in Ref. [43], used in GEMINI++ indicates a significant overestimation due to the large pairing correction calculation. Unfortunately, the available is absent for 11B∗, 15N∗ and their daughter nuclei after de-excitations. Therefore, we modify GEMINI++ through removing the back-shifted term in Eq. (2). In this case, the critical energy is closest to the separation energy. To validate the rationality of the removal of , we perform a simple calculation based on the formula in Sec. II 2B of Ref. [43] with the term containing removed. It is found that most of 11B∗, 15N∗ and their daughter nuclei can give a better agreement with the measured discrete levels due to removal of as discussed in the supplementary material. A similar calculation was initially performed to determine the boundary between discrete levels and continuum levels, which gives a rough result of 6 MeV. As a result, removing can enhance the and branching ratios, and correspondingly reduce the and emissions. For 11B∗, this modification will bring the predictions closer to the experimental observations. However, the predicted branching ratios of , and from 15N∗ de-excitations will deviate from their measured values.
@@
The daughter nuclei of de-excitations are usually left in discrete levels when they have a relatively low excitation energy. However, this situation does not appear in GEMINI++ since it only considers continuum levels. Here we calculate the influence of discrete levels, which can be implemented through modifying the GEMINI++ outputs on an event-by-event basis. Based on the estimated boundary of 6 MeV between discrete and continuum states, we only consider the discrete levels with the excitation energy less than about 6 MeV, and require that the decay information of every added discrete state is known clearly in the NNDC database [44]. Modifications of the GEMINI++ output events can be categorized into three types of situations for the excited daughter nuclei of 11B∗ and 15N∗:
If the excited energy of the decay nucleus in the last binary decay is higher than the highest discrete level considered obviously, this GEMINI++ event will be kept;
@@
Based on the above strategy, we modify the GEMINI++ outputs and find that including discrete levels can enhance the predicted branching ratios by about less than 20% for both ‘two-body decay‘ and ‘three-body decay’ in Fig. 1. This is because the aforementioned strategy can increase the kinetic energy of some of the finally emitted particles from below the detection threshold to above it. Note that the setting of discrete levels does not change the predicted relative branching ratios in Fig. 2.
@@ IV.2 GEMINI++4 results
Here we discuss the GEMINI++4 performance compared with the experimental data in Refs. [22, 23, 24]. Both and scenarios have removed the back-shifted term and added the discrete levels. The effects of step-by-step modifications can be found in the supplementary material. For the setting, the predicted branching ratios of 11B∗ de-excitations are in a good agreement with the measurements as shown in the left panel of Fig. 1. This is the first time that a statistical model code can simultaneously explain all four emissions of charged particles for both two-body and three-body decays of 11B∗, although there are still some minor differences. However, GEMINI++4 with can not account for the 15N∗ de-excitations well. It is clear that obvious differences in the right panel of Fig. 1 can be found for the , and decays. For the relative branching ratios in Fig. 2, GEMINI++4 with gives the better predictions than those from TALYS [6], NucDeEx [9] and ABLA [28].
@@ V Summary
In summary, we have investigated the de-excitations of highly excited 11B∗ and 15N∗ by use of the GEMINI++ code. Due to three issues of the critical excitation energy, discrete levels and suppression factors, GEMINI++ can not account for the experimental data well. After addressing the three issues, we develop a code of GEMINI++4 for neutrino experiments to handle de-excitaions of light nuclei based on the GEMINI++ code. For the settings of suppression factors, and have been assumed in this work. The predicted branching ratios of 11B∗ de-excitations are consistent with the measurements in the case. However, this setting can not explain the 15N∗ data. GEMINI++4 with can give the better agreement with the 11B∗ data, and partially account for the measured branching rations of 15N∗ de-excitations. Therefore, the setting is recommended for the relevant applications. This is the first time that a statistical model code can basically reproduce both 11B∗ and 15N∗ data. Moreover, the comparisons for each energy bin show that the agreement between the GEMINI++4 predictions and the experimental data is not coincidental. In the future, we plan to integrate GEMINI++4 into widely used neutrino generators, such as GENIE and NuWro.
@@ Acknowledgements.
- [13] R. Acciarri et al. [ArgoNeuT], Phys. Rev. D 99 (2019) no.1, 012002 doi:10.1103/PhysRevD.99.012002 [arXiv:1810.06502 [hep-ex]].
- [14] P. Abratenko et al. [MicroBooNE], Phys. Rev. D .109 (2024), 052007 doi:10.1103/PhysRevD.109.052007 [arXiv:2307.03102 [hep-ex]].
@@
- [17] W. Castiglioni, W. Foreman, I. Lepetic, B. R. Littlejohn, M. Malaker and A. Mastbaum, Phys. Rev. D 102 (2020) no.9, 092010 doi:10.1103/PhysRevD.102.092010 [arXiv:2006.14675 [physics.ins-det]].
@@
- [32] T. Golan, J. T. Sobczyk and J. Zmuda, Nucl. Phys. B Proc. Suppl. 229-232 (2012), 499-499 doi:10.1016/j.nuclphysbps.2012.09.136